OpenStudy (anonymous):

find the derivative of y=6xln(5x)

OpenStudy (anonymous):

product rule

OpenStudy (anonymous):

yeah looks like product rule,though i dont know the meaning of In?

OpenStudy (anonymous):

\[6\ln(5x)+6x/5x=6\ln(5x)+6/5\]

sam (.sam.):

\[dy/dx=6((\log5x)+1)\]